OpenStudy (anonymous):

write the expression in compact logarithm form: logx+ 1/2log(x+1)

OpenStudy (anonymous):

log x(x+1)^1/2

OpenStudy (math&ing001):

Hello @hill28 , To write the expression logx+ 1/2log(x+1) to compact logarithm form use the following rules : - a*log(x)=log(x^a) - log(x) + log(y) = log(xy) Tell me if you need more help.

OpenStudy (anonymous):

\[2\ln x-\ln (x+2)\] The directions say to write it as a simple logarithm. Just want to make sure I did it correctly.

OpenStudy (math&ing001):

@hill28 is that a second expression ? Because for the first, @beloab already pointed out the final answer (although I don't agree with giving answers directly).